OpenStudy (anonymous):

If the length of each side of a cube is tripled, is the surface area tripled? Use an example to support your answer.

OpenStudy (cwrw238):

surface area of a cube = 6*L^2 where L is the length of one side of the cube now work out the area if the length of a side is 3L

OpenStudy (anonymous):

can you explain more because i dont understand

OpenStudy (cwrw238):

each side of a cube is a square so its area is its length squared = L^2 now there are 6 sides on a cube so total surface area = 6* L^2

OpenStudy (anonymous):

OK that makes since

OpenStudy (cwrw238):

if length = 3L then total surface area = 6 * (3L)^2 = 6*9L^2 = 54L^2

OpenStudy (cwrw238):

so comparing the surface areas 54L^2 / 6 L^2 = 9 9 times bigger not tripled

OpenStudy (anonymous):

oh OK that makes much more since thanks you were very helpful for explaining

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

in fact as a general rule if you have two similar figures then ratio of their areas is L1 ^2 / L2 ^ 2 where L1 and L2 are corresponding lengths in the 2 figures

OpenStudy (cwrw238):

so in this case we can have length of 1 and 3 so ratio of ares is 1^2 : 3^2 = 1: 9

OpenStudy (anonymous):

wow you must be very smart to figure it out i suck at geometry

OpenStudy (cwrw238):

lol ty